find the next two terms 15,12,9,6,__,__
Answer by Edwin McCravy(20054)

To get from 15 to 12, you subtract 3, and if you do you get 15-3 = 12.
To get from 12 to 9, you subtract 3, and if you do you get 12-3 = 9.
To get from 9 to 6, you subtract 3, and if you do you get 9-3 = 6.
To get from 6 to __, you subtract 3, and if you do you get 6-3 = __.
To get from __ to __, you subtract 3, and if you do you get __-3 = __.


I'll bet you can now fill in those blanks.  If you can't then ask me
to help you in the thank-you note form below and I'll get back to you by
email.  Don't worry, I do this for fun and never charge a penny.

Edwin